# Combining inherent knowledge of vision-language models with unsupervised domain adaptation through self-knowledge distillation

### v2

@@ 3.4 Batch norm layer adjustment
The ResNet backbone employs batch norm layers. However, the distribution of the pretraining data is vastly different than the data used for the domain adaptation. The running average and mean change abruptly at the beginning of the adaptation process, shifting the data out of the learned distribution and resulting in failure. One solution would be to freeze the batch norm layers, but this can result in exploding feature norms, making the training unstable. In our work, we estimate the average and mean for the source and target data for each batch norm layer and adjust the learnable parameters $\beta$ and $\gamma$ so that the running mean and var equals that of our training data.
@@
where $\mu_{p}$ and $\sigma_{p}$ represent the running mean and variance of the pre-trained model, and $\mu_{c}$ and $\sigma_{c}$ represent the estimated mean and variance of the training data.
Table 3: Accuracy results on DomainNet dataset. ZS indicates zero-shot results of CLIP.
MCD [24] clp inf pnt qdr rel skt Avg CGDM [3] clp inf pnt qdr rel skt Avg MDD [37] clp inf pnt qdr rel skt Avg
clp - 15.4 25.5 3.3 44.6 31.2 24.0 clp - 16.9 35.3 10.8 53.5 36.9 30.7 clp - 20.5 40.7 6.2 52.5 42.1 32.4
inf 24.1 - 24.0 1.6 35.2 19.7 20.9 inf 27.8 - 28.2 4.4 48.2 22.5 26.2 inf 33.0 - 33.8 2.6 46.2 24.5 28.0
pnt 31.1 14.8 - 1.7 48.1 22.8 23.7 pnt 37.7 14.5 - 4.6 59.4 33.5 30.0 pnt 43.7 20.4 - 2.8 51.2 41.7 32.0
qdr 8.5 2.1 4.6 - 7.9 7.1 6.0 qdr 14.9 1.5 6.2 - 10.9 10.2 8.7 qdr 18.4 3 8.1 - 12.9 11.8 10.8
rel 39.4 17.8 41.2 1.5 - 25.2 25.0 rel 49.4 20.8 47.2 4.8 - 38.2 32.0 rel 52.8 21.6 47.8 4.2 - 41.2 33.5
skt 37.3 12.6 27.2 4.1 34.5 - 23.1 skt 50.1 16.5 43.7 11.1 55.6 - 35.4 skt 54.3 17.5 43.1 5.7 54.2 - 35
Avg 28.1 12.5 24.5 2.4 34.1 21.2 20.5 Avg 36 14 32.1 7.1 45.5 28.3 27.2 Avg 40.4 16.6 34.7 4.3 43.4 32.3 28.6
SCDA [13] clp inf pnt qdr rel skt Avg CDTrans [33] clp inf pnt qdr rel skt Avg SSRT [26] clp inf pnt qdr rel skt Avg
clp - 18.6 39.3 5.1 55 44.1 32.4 clp - 29.4 57.2 26 72.6 58.1 48.7 clp - 33.8 60.2 19.4 75.8 59.8 49.8
inf 29.6 - 34 1.4 46.3 25.4 27.3 inf 57 - 54.4 12.8 69.5 48.4 48.4 inf 55.5 - 54 9 68.2 44.7 46.3
pnt 44.1 19 - 2.6 56.2 42 32.8 pnt 62.9 27.4 - 15.8 72.1 53.9 46.4 pnt 61.7 28.5 - 8.4 71.4 55.2 45
qdr 30.0 4.9 15 - 25.4 19.8 19 qdr 44.6 8.9 29 - 42.6 28.5 30.7 qdr 42.5 8.8 24.2 - 37.6 33.6 29.3
rel 54 22.5 51.9 2.3 - 42.5 34.6 rel 66.2 31 61.5 16.2 - 52.9 45.6 rel 69.9 37.1 66 10.1 - 58.9 48.4
skt 55.6 18.5 44.7 6.4 53.2 - 35.7 skt 69 29.6 59 27.2 72.5 - 51.5 skt 70.6 32.8 62.2 21.7 73.2 - 52.1
Avg 42.6 16.7 37 3.6 47.2 34.8 30.3 Avg 59.9 25.3 52.2 19.6 65.9 48.4 45.2 Avg 60 28.2 53.3 13.7 65.3 50.4 45.2
PMTrans [40] clp inf pnt qdr rel skt Avg ZS [22] clp inf pnt qdr rel skt Avg SKD (Ours) clp inf pnt qdr rel skt Avg
clp - 34.2 62.7 32.5 79.3 63.7 54.5 clp - 48.3 68.5 14.0 84.6 65.8 56.3 clp - 52.7 72.0 36.3 85.1 71.5 63.5
inf 67.4 - 61.1 22.2 78 57.6 57.3 inf 73.1 - 68.5 14.0 84.6 65.8 61.2 inf 78.5 - 72.2 30.4 85.0 69.5 67.1
pnt 69.7 33.5 - 23.9 79.8 61.2 53.6 pnt 73.1 48.3 - 14.0 84.6 65.8 57.2 pnt 78.7 49.8 - 29.6 84.5 69.7 62.4
qdr 54.6 17.4 38.9 - 49.5 41 40.3 qdr 73.1 48.3 68.5 - 84.6 65.8 68.1 qdr 70.7 22.9 50.9 - 82.4 62.3 57.8
rel 74.1 35.3 70 25.4 - 61.1 53.2 rel 73.1 48.3 68.5 14.0 - 65.8 54.0 rel 80.3 53.5 74.2 30.2 - 71.1 61.9
skt 73.8 33.0 62.6 30.9 77.5 - 55.6 skt 73.1 48.3 68.5 14.0 84.6 - 57.7 skt 81.0 52.0 73.0 36.6 84.6 - 65.4
Avg 67.9 30.7 59.1 27 72.8 56.9 52.4 Avg 73.1 48.3 68.5 14.0 84.6 65.8 59.1 Avg 77.9 46.2 68.5 32.6 84.3 68.8 63.0
@@ 4 Experiments
We evaluate our proposed method on three different domain adaptation benchmarks, Office-Home, VisDA, and DomainNet. We show that we can improve the baselines significantly. In ablation studies, we further investigate the contribution of the different parts of our proposed algorithm.
4.1 Setup
@@
Implementation details: We built up our implementation on the CDAN implementation of [15]. We do all experiments with the CLIP-trained versions of the ResNet backbone and transformer backbone. For the ResNet backbone, we follow most other publications and employ ResNet-50 for OfficeHome and ResNet-101 for VisDA. For the transformer backbone, we employ ViT-B/16. We employ the seven ImageNet templates subset published on the CLIP Git Hub for the language encoder to estimate the zero-shot predictions. For OfficeHome we train each run for $100$ episodes, for VisDA and DomainNet we train for $50$ episodes. For OfficeHome and DomainNet, we employ AdamW optimizer with a learning rate of $5e^{-6}$ for the backbone and $5e^{-5}$ for the newly initialized layers. For VisDA we employ a lower learning rate of $1e^{-6}$, and $1e^{-5}$ respectively. We employ the GSDE strategy for a total of $N=5$ runs. Each experiment is run for three different seeds, except for the DomainNet which we only train for one seed.
@@
Table 4: Contributions of different parts of our algorithm. Aug means using strongly augmented images alongside the weakly augmented ones, while GSDE uses only the prediction scores of the previous run, GSDE+ adds zero-shot prediction scores according to eq. 5.
ResNet ViT
$L_{AD}$ $L_{KD}$ Aug GSDE GSDE+ VisDA OH VisDA OH
$\checkmark$ 72.73 65.99 74.25 81.74
$\checkmark$ 87.65 69.06 90.28 83.27
$\checkmark$ $\checkmark$ 88.21 73.43 90.44 87.66
$\checkmark$ $\checkmark$ $\checkmark$ 88.62 75.48 90.37 87.97
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 88.90 78.72 91.22 89.08
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 89.09 78.78 91.20 89.24
Table 5: Generalization to unseen data.
train test comb
ZS 59.05 59.11 59.07
Ours 63.32 62.38 63.04
@@ 4.2 Results
Results for Office-Home: The results for the Office-Home dataset are shown in Tab. 1. The zero-shot predictions already outperform some of the earlier works but are considerably below current state-of-the-art algorithms. Using our method, we improve the average accuracy by $7.3$pp for the ResNet-50 backbone and $7.0$pp for the ViT backbone. An improvement can be seen for all $12$ adaptation tasks. For the ResNet backbone, we improve the state-of-the-art by around $4.3$pp.
Results for VisDA-2017: The results for the VisDA-2017 dataset are shown in Tab. 2. Again, the zero-shot predictions already outperform some of the earlier works but are considerably below current state-of-the-art algorithms. Using our method, we improve the average accuracy by $4.4$pp for the ResNet-101 backbone and $2.6$pp for the ViT backbone, letting us outperform current state-of-the-art methods.
Results for DomainNet: The results for the DomainNet dataset are shown in Tab. 3. The zero-shot predictions already outperform current state-of-the-art methods considerably. By using our method we further increase the average accuracy by $3.9$pp. We achieve the biggest increase of performance over the zero-shot predictions for the target set of quickdraw, increasing the accuracy by $18.6$pp. However, when using the quickdraw domain as source data, the accuracy decreases considerably by around $10.3$pp compared to the zero-shot predictions. The quickdraw domain has a large domain gap to the other domains. This large gap introduces a negative transfer resulting in a drop in accuracy.
@@ 5 Ablation studies
Parameter for adjusting zero-shot distribution $\tau$: In this part, we evaluate the effect of the parameter $\tau$. Figure 4 shows the accuracy for different values of $\tau$. None means that no distribution shift was applied to the zero-shot predictions. It can be seen that applying the distribution shift is essential for the method. Employing a shift of $\tau=0.9$ increases the accuracy by $10.9$pp for ResNet and $6.2$pp for ViT backbone. For the ViT backbone, the accuracy increases constantly until a value of $0.9$, for which the accuracy peaks. For the ResNet backbone, it seems that even higher values of $\tau$ yield better results. In all our experiments, we employ a value for $\tau$ of $0.9$.
Contribution of losses: In this part, we evaluate the contribution of the different losses to our method. The combination of $L_{AD}$ and $L_{KD}$ improves the accuracy in all of the four settings. Especially for the OfficeHome dataset, where adversarial domain adaptation and knowledge distillation achieve similar results, the boost is the largest. Data augmentation further slightly increases the accuracy. And we achieve another boost by introducing GSDE. Introducing the zero-shot predictions into the GSDE algorithm increases the accuracy further a slight bit.
Generalization to unseen data: Since we only used the training sets for the adaptation task in the DomainNet dataset, it allows us to evaluate the generalization ability to unseen data. The results are shown in Tab. 5. As expected the zero-shot accuracy is similar between the train and test set. Our method has a slight drop of around $1$pp for the test set, meaning that there is a slight overfitting to the seen data. However it still significantly outperforms the zero-shot prediction accuracy. Since the train dataset is much larger than the test dataset, the combined results are closer to the train dataset.
T-SNE visualization: We visualize the feature embedding using T-SNE for the addition of the different losses as well as for the CLIP pre-trained model in 5. Since the CLIP model is not trained classification but aligning visual and language features, the distribution is very spread out (see Fig. 4(a)). Interestingly the source and target domain are quite separated. Using the KD loss (and classification loss) aligns the different classes into clusters, but does not do a good job of aligning the features (see Fig. 4(c)). The CDAN loss aligns both domains well but suffers from misalignment between different classes (see Fig. 4(e)). By combining the two methods, we get well-aligned domains with fewer misalignments (see Fig. 4(g)).
(a) CLIP.
(b) CLIP.
(c) KD.
(d) KD.
(e) CDAN.
(f) CDAN.
(g) Combined.
(h) Combined.
[…]
6 Discussion
In this work, we presented a novel way to combine the knowledge of vision-language models with knowledge gained from a source dataset through unsupervised domain adaptation. We maintain the inherent knowledge of the vision-language model via a knowledge distillation loss. To effectively distill the knowledge we introduced a distribution adjustment for the prediction probabilities. For the domain adaptation loss, we employed CDAN as it is effective for both CNN and transformer-based backbones. We showed the effectiveness of our method for both backbones on three different datasets, outperforming current state-of-the-art methods. We showed that combining unsupervised domain adaptation with the inherent knowledge of vision-language models improves the performance significantly over using only one of the knowledge sources. It should be stated, that it can be expected to further increase the performance by choosing a stronger method for the domain adaptation task.
@@ Acknowledgment
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015, JSPS KAKENHI Grant Number JP21K17799, and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo.

### v4

@@ 3.4.1 Batch norm layer adjustment
The ResNet backbone employs batch norm layers. However, the distribution of the pretraining data is vastly different than the data used for the domain adaptation. To overcome this problem we employ domain-specific BN layers as in [1] and adjust the BN layers to the statistics of the data of their respective domain similar to [15]. We estimate the average and mean for the source and target data for each batch norm layer and adjust the learnable parameters $\beta$ and $\gamma$ so that the running mean and var equals that of our training data (one adjusted to source domain and one to the target domain).
@@
where $\mu_{p}$ and $\sigma_{p}$ represent the running mean and variance of the pre-trained model, and $\mu_{c}$ and $\sigma_{c}$ represent the estimated mean and variance of the training data. We want to highlight that this is only done for the ResNet backbone since the ViT backbone employs Layer Normalization instead of BN.
3.4.2 Zero-shot predictions
To increase the quality of zero-shot predictions, we try to keep the original ratios of the images. The smaller side of an image gets resized to 224 pixels and the larger side gets resized by the same factor but rounded to the closes multiple of 16 - according to the patchsize of the ViT-backbone (32 for ResNet-backbone due to the attention pooling). We interpolate the positional embedding in order for the network to process the different image sizes.
3.4.3 DAPL
Domain Adaptation via Prompt Learning (DAPL) [7] focuses on adapting the prompts of the text encoder. It freezes both text and visual encoder during the training. In contrast to this, our method focuses on adapting the visual encoder. We combine these two methods in a sequential manner. In a first step, we run the DAPL algorithm and use the adapted text prompts to estimate the probabilities for the source and target data, which are then employed instead of the zero-shot predictions.
@@ 4 Experiments
We evaluate our proposed method on three different domain adaptation benchmarks, Office-Home, VisDA, and DomainNet. We show that we can improve the baselines significantly.
Figure 3: Accuracy for adaptation of OfficeHome dataset for different source domain expansion percentages. The green line employs only the strong guidance, while the orange line employs both guidance. Additionally, the baselines of only using the adversarial domain adaptation (CDAN) and the CLIP zero-shot accuracy (ZS) are plotted.
4.1 Experiment settings
@@
Implementation details: We built up our implementation on the CDAN implementation of [17]. We do all experiments with the CLIP-trained versions of the ResNet backbone and transformer backbone. For the ResNet backbone, we follow most other publications and employ ResNet-50 for OfficeHome and ResNet-101 for VisDA. For the transformer backbone, we employ ViT-B/16. We employ the seven ImageNet templates subset of the CLIP Git Hub page as text prompts for estimating the zero-shot predictions. For OfficeHome we train each run for $100$ episodes, for VisDA and DomainNet we train for $50$ episodes. For OfficeHome and DomainNet, we employ AdamW optimizer with a learning rate of $5e^{-6}$ for the backbone and $5e^{-5}$ for the newly initialized layers. For VisDA we employ a lower learning rate of $1e^{-6}$, and $1e^{-5}$ respectively. We employ a batch size of $64$. We load each image once with weak and once with strong data augmentation, but don’t discriminate between the augmentations in the processing pipeline. We run two different versions of our method. The first version (V1) trains with a source expansion set to $50\%$. The second version (V2) trains at first with a source expansion set to $33.3\%$ in a first run, and $66.6\%$ in a second run, where the first run employs the zero-shot predictions and the second a mixture between zero-shot and previous run predictions for the source expansion. Due to the two training runs, the second version requires about double the training time. We employ the official implementation of DAPL for the respective experiments.
@@ 4.2 Results
Results for Office-Home: The results for the Office-Home dataset are shown in Tab. 1. All versions of our method vastly outperform earlier work. The addition of DAPL improves the results by around $3-4\%$pts in the case of the ResNet backbone and around $2\%$pts for the ViT backbone. The second version using an iterative training scheme with 2 runs increases the accuracy in all cases, this comes however at the cost of roughly double the training time.
Results for VisDA-2017: The results for the VisDA-2017 dataset are shown in Tab. 2. Our method just by itself already outperforms most of the earlier works. Our method slightly performs worse than EUDA only for the case of the ViT-backbone and without using DAPL. However, once we add DAPL, we significantly outperform the next best-performing method for the ResNet and ViT backbone.
Results for DomainNet: The results for the DomainNet dataset are shown in Tab. 3. The CLIP-based adaptation methods clearly outperform conventional methods by upwards of $10\%$pts. Our method significantly outperforms current state-of-the-art methods, increasing the average accuracy by $1.4\%$pts or $2.4\%$pts respectively over the next best performing method. Version 2, the iterative version of our method increases the accuracy by $1\%$pts over version 1. This however comes again at the price of the increase in training time.
@@ 5 Ablation studies
Table 4: Effect of batch-norm layer adjustment. BN layer non adjusted (C), BN layer adjusted for source data (C-SAd), adjusted for target data (C-TAd), domain specific BN non adjusted (D), and domain specific layer adjusted to their specific domain D-STAd.
Method S S-SAd S-TAd D D-STAd
OH 78.5 78.8 78.9 78.5 79.7
Combination of strong and weak guidance The effect of combining the strong and weak guidance can be seen in Fig. 3. It should be noted in the case of $0\%$ source expansion, the strong guidance is equivalent to CDAN, and the strong-weak guidance is equivalent to only weak guidance. The guidance vastly improves over the baselines, and the combination of both even further improves the results. According to the results we chose a source expansion of $50\%$ in the single run experiments of our evaluations.
@@
Parameter for shifting the zero-shot predictions $\tau$: The effect of the parameter $\tau$ is shown in Fig. 4. None means that no adjustment was applied to the zero-shot predictions. It can be seen that applying the ZS probably adjustment is essential for the method. Employing a target mean winning probability of $\tau=0.9$ increases the accuracy by $10.9$pp for ResNet and $6.2$pp for ViT backbone. For the ViT backbone, the accuracy increases constantly until a value of $0.9$, for which the accuracy peaks. For the ResNet backbone, it seems that even higher values of $\tau$ yield better results. In all our experiments, we employ a value for $\tau$ of $0.9$.
Batchnorm layer adjustment: In Tab. 4 it can be seen that adjusting the BN layer to either source or target gains a small increase in performance. Changing the BN to domain specific BN does not improve the performance just by itself, the performance increases just in combination with adjusting each BN layer to its domain. The performance gain of $1.2\%$pts is quite considerable. But even without adjusting the BN layer, we still outperform all other methods.
Combining the weak-strong guidance with other DA methods: In this part, we evaluated how well the weak-strong guidance fares other DA methods. We did not employ BN adjustment for this experiment. The results are shown in Tab. 5. While MDD outperforms CDAN for the ResNet backbone, it underperforms for the transformer backbone. For ViT, choosing the transformer-based method SSRT performs best. CDAN performs good for both backbones.
Influence of quality of ZS predictions: In this part, we evaluated the influence of the quality of ZS predictions on the adaptation. We generated different ZS predictions with a combination of different text templates (single vs multiple) and image size (rectangular vs ratio preserving). Single template is ”a + cls” while multiple is the ImageNetSmall list from CLIP. Rectangular image size resizes the image to 224x224, while ratio preserving resizes the smaller side to 224 and the larger side to its ratio preserving size (it uses adaptive input size). It can be seen (Tab. 6) that the quality of the zero-shot prediction has a big influence. However, while the difference between best and worst ZS prediction is $5\%$pts (ResNet); $3.8\%$pts (ViT), it’s only $2.4\%$pts; $1\%$pts after the adaption.
Table 5: Combination of weak-strong guidance with other DA methods. The results shown are for the OfficeHome dataset. Besides CDAN, we employed MDD, SENTRY, and SSRT (transformer-based method).
Method CDAN MDD SENTRY SSRT
ResNet 78.5 80.0 78.2 -
ViT 89.7 87.9 85.7 90.1
Table 6: Effect of ZS prediction quality on adaptation. The ZS predictions are generated with different combinations of text templates (Single vs Multiple) and image size (Rectangular vs ratio Preserving).
S-R M-R S-P M-P
ResNet ZS 70.3 73.5 73.0 75.3
Results 77.3 78.4 78.8 79.7
ViT ZS 80.4 83.2 83.2 84.2
Results 88.7 89.3 89.1 89.7
6 Conclusion
In this work, we presented a novel way to combine the knowledge of vision-language models with knowledge gained from a source dataset through a strong-weak guidance scheme. For the strong guidance, we expand the source dataset with the most confident samples of the target dataset. Additionally, we employ a knowledge distillation loss as weak guidance. The strong guidance uses hard labels but is only applied to the most confident predictions from the target dataset. Conversely, the weak guidance is employed to the whole dataset but uses soft labels. For the domain adaptation loss, we employed CDAN as it is effective for both CNN and ViT-based backbones, but also show that other methods can be used. Furthermore, we showed that our method is complementary to prompt learning methods and a combination can further improve the accuracy.
@@ Acknowledgment
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015 and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo.